OpenStudy (anonymous):

Solve using your choice of methods. x² + 7x + 10 = 0

OpenStudy (anonymous):

I would complete the square which is the same as factoring it out.

OpenStudy (anonymous):

\[(x+a)(x+b)=0 \\ a+b=7 \\a*b=10\] Where a and b are two numbers that when you add them together you will get 7 and if you multiply them together you will get 10

OpenStudy (anonymous):

That is shown in the system of equation that I posted. You can solve the systems and see what a and b are.

OpenStudy (anonymous):

can you tell me what a and b are?

OpenStudy (anonymous):

(x+5)(x+2)

OpenStudy (anonymous):

a=1,b=7,c=10

OpenStudy (anonymous):

You really want to use the quadratic formula. I'm trying to show you a different way to solve this.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

show me a different way.

OpenStudy (anonymous):

@sarah43 is in the right direction. Can you show jeletta1 how to do it?

OpenStudy (anonymous):

ok. look at it like this x² + 7x + 10 = 0 you have to find 2 numbers that when MULTIPLIED equal 10 the same 2 numbers should equal 7 when ADDED so lets look at the factors of 10. 5*2=10! and 5+2=7! so (x+5)(X+2) is the answer because when you open it up u will get x² + 7x + 10 = 0